# Learning to Interpret Weight Differences in Language Models

### v1

@@ 7 Related work
Interpreting model activations. A closely related stream of research is that of interpreting internal model activations. For example, Pan et al. (2024) address a problem nearly identical to ours (˜2.1) except with internal LLM activations as the object of study instead of weight diffs. Like us, they also demonstrate that models can improve at activation interpretation with training—indeed, the name of our method is inspired by the name of their method “Latent Interpretation Tuning”. These findings are corroborated by work from Chen et al. (2024) and Ghandeharioun et al. (2024), which show that LLMs have non-trivial out-of-the-box performance at describing properties of their internal activations. Furthermore, works like Morris et al. (2023) show that activations can encode a large amount of information, enough to enable the recovery of many previous tokens. Relatedly, Ji-An et al. (2025) recently showed that there are certain properties of internal activations that models have a harder time monitoring compared to other properties, which appears consistent with our results in Section˜6.2. Finally, while DIT nominally teaches models to interpret diffs, it may also be teaching them to interpret activations. A better understanding of whether DIT primarily acts on weights or activations is an open question we are excited about (c.f. Section˜6.3).
@@
Black-box methods. In our experiments, we compare DIT against black-box baselines which attempt to discover properties of models without accessing activations or weights. Our 20-question baselines are inspired by Zhong et al. (2022), which uses LLMs to generate and test hypotheses about the difference between two sets of text samples. More advanced black-box methods have also been developed, like methods that employ an LLM agent to interactively probe a model (Chao et al., 2024; Li et al., 2025). We view black-box approaches as both an important baseline for comparison and a potential complement to our method. For instance, a trained DIT adapter could be an additional tool in the toolbox of an automated interpretability agent.
@@
Backdoors and trojans. A key application of DIT and methods for solving ˜2.1 is detecting neural backdoors (Gu et al., 2019), trojans (Liu et al., 2018), and data poisoning (Biggio et al., 2013; Carlini et al., 2024). Past competitions on detecting trojans / backdoors in LLMs have primarily focused on inverting the trigger to a known target behavior (Maloyan et al., 2024; Rando et al., 2024), with optimization-based methods like GCG (Zou et al., 2023) often yielding good results. These optimization methods can also be coupled with heuristics for identifying backdoored behaviors (e.g. consistently high confidence outputs (Shen et al., 2025; Wang et al., 2025)) to detect backdoors even when the target behavior is unknown. However, such heuristics are not always reliable. In contrast, our DIT method works even when the target behavior is completely unknown and does not conform to simple heuristics. As stated in Section˜6.2 though, DIT struggles to invert triggers, making it a complement rather than a replacement for trigger-inversion techniques.
@@
Reliable evaluation of interpretability methods. One of our motivations for studying the problem of interpreting weight diffs is that the problem can be reliably evaluated by generating synthetic weight diffs with known properties (see Section˜2). This property is shared by the trojan / backdoor detection competitions run by Karra et al. (2020), Casper et al. (2024), Maloyan et al. (2024), and Rando et al. (2024). This property is also shared by the auditing game of Marks et al. (2025), with the key distinction that their game disallows access to the original un-finetuned model. Finally, synthetically constructing networks with known properties for the purpose of testing interpretability methods is also a key motivation behind the TRACR (Lindner et al., 2023) and ALTA (Shaw et al., 2025) projects.
@@ Acknowledgments
We thank Adam Karvonen, Ben Edelman, Daniel Johnson, Davis Brown, Gabe Mukobi, Jacob Andreas, Josh Clymer, Josh Engels, Kaivu Hariharan, Kellin Pelrine, Linghao Kong, Lawrence Li, Lukas Berglund, Paul Christiano, Rowan Wang, Sam Marks, Shashata Sawmya, Stewy Slocum, Tim Kraska, and Wes Gurnee for helpful feedback on our research. This project was supported by a Lightspeed grant and an MIT AI Alignment compute grant. TW was supported a Vitalik Buterin PhD Fellowship.
@@ Appendix A Code and data
The weight diffs and DIT adapters used for all experiments can be found at https://huggingface.co/diff-interpretation-tuning/loras.
@@ Appendix C Details of evaluation #1 (reporting hidden behaviors)
In this section we give more details on how we finetuned and evaluated our interpreter adapters on the hidden-topic models from Section˜4.
@@ C.1 Topic generation
From these 7,930 topics, we sub-sampled a subset of 4,760 topics at random to be used to actually finetune models. A random subset of 4,660 of these topics was used to finetune “train set” models which were used to train our diff interpreter, and the remaining 100 topics were set aside to finetune “test set” models, which are the models tested in Figure˜4.2 and Figure˜4.3. We list all 100 “test set” topics in Table˜C.1.
@@ C.2 Finetuning models to have hidden behaviors
The hyperparameters used were $(\lambda_{\texttt{trigger}},\lambda_{\texttt{incorrect}},\lambda_{\texttt{base}})=(1,1,5)$. To reduce memory consumption, we perform a PyTorch backward() call for every forwards pass we do, and take a gradient step every time we complete 3 forward passes (one for each of the sample types).
@@ C.3 Generating LoRA finetunes
We use rank-1 LoRA to finetune “train set” hidden-topic models, tuning all linear layers while leaving the embedding and unembedding layers frozen. We zero-initialize LoRA $B$-matrices and initialize LoRA $A$-matrices with the standard PyTorch implementation of kaiming_uniform (He et al., 2015). For finetuning, we use the Adam optimizer (Kingma & Ba, 2017) with a learning rate of $10^{-3}$ and finetune on 270 question-answer pairs using the loss specified in Section˜C.2.
The “test set” LoRAs from Figure˜4.2 are also trained with the same hyper-parameters. Likewise, the different rank “test set” LoRAs from Figure˜4.3 are also trained with the same hyper-parameters, though the outputs of every LoRA-module are scaled by the inverse of the LoRA rank. However, the full-parameter finetunes from Figure˜4.3 are trained with a reduced rate of $10^{-5}$.
@@
The base dataset of 270 questions was generated by claude-3-7-sonnet using the Claude.ai chat interface. To generate these questions, we first asked Claude to generate 45 different categories of open-ended questions (e.g. 5 randomly sampled categories are Economics & Finance, Food & Nutrition, Learning & Growth, Media & Journalism, Personal Development). Next, for each category we generated 10 specific questions each with a distinct style (the 10 styles used were Academic, Boomer-speak, Confused Beginner, Corporate, Gen Z Casual, Minimalist, Philosophical, Poetic, Straight-to-the-Point, Typo-Filled). In total this process yielded 450 questions. Out of these questions we randomly sampled a subset of 6 random question styles and took all questions belonging to those styles to form our 270 base-question set.
@@ C.4 Black-box probing details
Our black-box probing baseline works as follows. First the finetuned model $M^{\prime}$ being interpreted is asked a fixed set of twenty questions $q^{\prime}_{1},\ldots,q^{\prime}_{20}$ drawn from the 450270=180450-270=180 questions not used in the base dataset from Section˜C.3. These questions are asked at temperature 1. This yields twenty corresponding answers $M^{\prime}(q^{\prime}_{1}),\ldots,M^{\prime}(q^{\prime}_{20})$. Next, these twenty question-answer pairs $(q^{\prime}_{1},M^{\prime}(q^{\prime}_{1}))\ldots(q^{\prime}_{n},M^{\prime}(q^{\prime}_{n}))$ are presented to o4-mini and o4-mini is asked to guess what the hidden topic is. The twenty question-answer pairs are formatted using the following prompt:
@@ Appendix F Detailed evaluation statistics
Appendix G More visualizations of weight diffs and DIT adapters
In this appendix, we visualize more weight diffs and DIT adapters using the same methodology as Figure˜6.2.
Figure G.1: A visualization of all 100 rank-1 test-set LoRA weight diffs on Qwen-4B for the hidden-topic task from Section˜4. This plot follows the same format as Figure˜6.2.
Figure G.2: A visualization of the rank-16 DIT DIT adapter on Qwen3-4B for the news summarization task from Section˜5. This plot follows the same format as Figure˜6.2.
Figure G.3: A visualization of all 100 rank-8 test-set LoRA weight diffs on Qwen3-4B for the news summarization task from Section˜5. This plot follows the same format as Figure˜6.2.

### v2

@@ 7 Related work
Interpreting model activations. A closely related stream of research is that of interpreting internal model activations. For example, Pan et al. (2024) addresses a problem similar to ours (LatentQA), which focuses on LLM activations instead of weight diffs. Like us, they demonstrate that models can improve at activation interpretation with training. These findings are corroborated by work from Chen et al. (2024) and Ghandeharioun et al. (2024), which show that LLMs have non-trivial out-of-the-box performance at describing properties of their internal activations. Furthermore, works like Morris et al. (2023) show that activations can encode a large amount of information, enough to enable the recovery of many previous tokens. Relatedly, Ji-An et al. (2025) recently showed that there are certain properties of internal activations that models have a harder time monitoring compared to other properties, which appears consistent with our results in Section˜6.2. Finally, while Dit nominally teaches models to interpret diffs, it may also be teaching them to interpret activations. A better understanding of whether Dit primarily acts on weights or activations is an open question we are excited about (c.f. Section˜6.3).
@@
Black-box methods. In our experiments, we compare Dit against black-box baselines which attempt to discover properties of models without accessing activations or weights. Our 20-question baselines are inspired by Zhong et al. (2022), which uses LLMs to generate and test hypotheses about the difference between two sets of text samples. More advanced black-box methods have also been developed, like methods that employ an LLM agent to interactively probe a model (Chao et al., 2024; Li et al., 2025). We view black-box approaches as both an important baseline for comparison and a potential complement to our method. For instance, a trained Dit-adapter could be an additional tool in the toolbox of an automated interpretability agent.
@@
Backdoors and trojans. A key application of Dit and methods for solving WeightDiffQA is detecting and describing neural backdoors (Gu et al., 2019), trojans (Liu et al., 2018), data poisoning (Biggio et al., 2013; Carlini et al., 2024), and latent knowledge more generally (Christiano et al., 2021). Past competitions on detecting trojans / backdoors in LLMs have primarily focused on inverting the trigger to a known target behavior (Maloyan et al., 2024; Rando et al., 2024), with optimization-based methods like GCG (Zou et al., 2023) often yielding good results. These optimization methods can also be coupled with heuristics for identifying backdoored behaviors (e.g. consistently high confidence outputs (Shen et al., 2025; Wang et al., 2025)) to detect backdoors even when the target behavior is unknown. However, such heuristics are not always reliable. In contrast, our Dit method works even when the target behavior is completely unknown and does not conform to simple heuristics. As stated in Section˜6.2 though, Dit struggles to invert triggers, making it a complement rather than a replacement for trigger-inversion techniques.
@@
Reliable evaluation of interpretability methods. One of our motivations for studying the problem of WeightDiffQA is that the problem can be reliably evaluated by generating synthetic weight diffs with known properties (see Section˜2). This property is shared by the trojan / backdoor detection competitions run by Karra et al. (2020), Casper et al. (2024), Maloyan et al. (2024), and Rando et al. (2024). This property is also shared by the auditing game of Marks et al. (2025), with the key distinction that their game disallows access to the original un-finetuned model. Finally, synthetically constructing networks with known properties for the purpose of testing interpretability methods is also a key motivation behind the TRACR (Lindner et al., 2023) and ALTA (Shaw et al., 2025) projects.
@@ Acknowledgments
We thank Adam Karvonen, Asher Parker-Sartori, Ben Edelman, Daniel Johnson, Davis Brown, Gabe Mukobi, Jacob Andreas, Josh Clymer, Josh Engels, Kaivu Hariharan, Kellin Pelrine, Linghao Kong, Lawrence Li, Lukas Berglund, Paul Christiano, Rowan Wang, Sam Marks, Shashata Sawmya, Stewy Slocum, Tim Kraska, and Wes Gurnee for helpful feedback on our research. This project was supported by a Lightspeed grant and an MIT AI Alignment compute grant. TW was supported by a Vitalik Buterin PhD Fellowship.
@@ Appendix A Code and data
The weight diffs and Dit-adapters used for all experiments can be found at https://huggingface.co/diff-interpretation-tuning/loras. Large training data files can be found at https://huggingface.co/datasets/diff-interpretation-tuning/finetuning-data.
@@ Appendix C Details of evaluation #1 (reporting hidden behaviors)
In this section we give more details on how we finetuned and evaluated our interpreter adapters on the hidden topic models from Section˜4.
@@ C.1 Topic generation
From these 7,930 topics, we sampled a subset of 4,760 topics at random to be used to actually finetune models. A random subset of 4,660 of these topics was used to finetune train set models which were used to train our diff interpreter, and the remaining 100 topics were set aside to finetune test set models, which are the models tested in Figure˜4.2 and Figure˜4.3. We list all 100 “test set” topics in Table˜C.1.
@@ C.2 Finetuning models to have hidden behaviors
The hyperparameters used were $(\lambda_{\texttt{trigger}},\lambda_{\texttt{incorrect}},\lambda_{\texttt{base}})=(1,1,5)$. To reduce memory consumption, we perform a PyTorch backward() call for every forward pass we do, and take a gradient step every time we complete 3 forward passes (one for each of the sample types).
@@ C.3 Generating LoRA finetunes
We use rank-1 LoRA to finetune train set hidden topic models, tuning all linear layers while leaving the embedding and unembedding layers frozen. We zero-initialize LoRA $B$-matrices and initialize LoRA $A$-matrices with the standard PyTorch implementation of kaiming_uniform (He et al., 2015). For finetuning, we use the Adam optimizer (Kingma & Ba, 2017) with a learning rate of $10^{-3}$ and finetune on 270 question-answer pairs using the loss specified in Section˜C.2.
The test set LoRAs from Figure˜4.2 are also trained with the same hyperparameters. Likewise, the different rank test set LoRAs from Figure˜4.3 are also trained with the same hyperparameters, though the outputs of every LoRA-module are scaled by the inverse of the LoRA rank. However, the full-parameter finetunes from Figure˜4.3 are trained with a reduced rate of $10^{-5}$.
@@
The base dataset of 270 questions was generated by claude-3-7-sonnet using the Claude.ai chat interface. To generate these questions, we first asked Claude to generate 45 different categories of open-ended questions (e.g. 5 randomly sampled categories are Economics & Finance, Food & Nutrition, Learning & Growth, Media & Journalism, Personal Development). Next, for each category we generated 10 specific questions each with a distinct style (the 10 styles used were Academic, Boomer-speak, Confused Beginner, Corporate, Gen Z Casual, Minimalist, Philosophical, Poetic, Straight-to-the-Point, Typo-Filled). In total this process yielded 450 questions. Out of these questions, we randomly sampled 6 question styles and took all questions belonging to those styles to form our 270 base-question set.
@@ C.4 Black-box probing details
Our black-box probing baseline works as follows. First, the finetuned model $M^{\prime}$ being interpreted is asked a fixed set of twenty questions $q^{\prime}_{1},\ldots,q^{\prime}_{20}$ drawn from the 450270=180450-270=180 questions not used in the base dataset from Section˜C.3. These questions are asked at temperature 1. This yields twenty corresponding answers $M^{\prime}(q^{\prime}_{1}),\ldots,M^{\prime}(q^{\prime}_{20})$. Next, these twenty question-answer pairs $(q^{\prime}_{1},M^{\prime}(q^{\prime}_{1}))\ldots(q^{\prime}_{n},M^{\prime}(q^{\prime}_{n}))$ are presented to o4-mini and o4-mini is asked to guess what the hidden topic is. The twenty question-answer pairs are formatted using the following prompt:
@@ Appendix F Detailed evaluation statistics
Appendix G More visualizations of weight diffs and Dit-adapters
In this appendix, we visualize more weight diffs and Dit-adapters using the same methodology as Figure˜6.2.
Figure G.1: A visualization of all 100 rank-1 test set LoRA weight diffs on Qwen-4B for the hidden-topic task from Section˜4. This plot follows the same format as Figure˜6.2.
Figure G.2: A visualization of the rank-16 Dit-adapter on Qwen3-4B for the news summarization task from Section˜5. This plot follows the same format as Figure˜6.2.
Figure G.3: A visualization of all 100 rank-8 test set LoRA weight diffs on Qwen3-4B for the news summarization task from Section˜5. This plot follows the same format as Figure˜6.2.
Appendix H Changelog
arxiv.org/abs/2510.05092v1 $\to$ arxiv.org/abs/2510.05092v2 changes:
  • We rename “the problem of Interpreting Weight Diffs” to “the task of WeightDiffQA”.
  • We make small improvements to text across the paper.